OpenStudy (anonymous):

Which of the following functions have a domain of (-infinity, infinity), and a range of (2, infinity)? Check all that apply A. f(x)= 3^x+2 B. f(x_= 3^x-2 C. f(x)= 0.25^x +2 D. f(x)= 0.25^x-2

jimthompson5910 (jim_thompson5910):

how far did you get here?

OpenStudy (anonymous):

well honestly i have a hard time with problems like these i always guess could you explain how to start ?

jimthompson5910 (jim_thompson5910):

what is the domain in general? What does the term mean?

jimthompson5910 (jim_thompson5910):

If you had to define it in your own words, what would you say?

OpenStudy (anonymous):

domain means all the x terms or the independent variable

jimthompson5910 (jim_thompson5910):

basically all of the allowed inputs of a function

jimthompson5910 (jim_thompson5910):

for something like f(x)= 3^x+2, are there any restrictions on x?

OpenStudy (anonymous):

no i do not think so

jimthompson5910 (jim_thompson5910):

you are correct, there are no restrictions you can plug in ANY real number for x and get some output

jimthompson5910 (jim_thompson5910):

so that's why the domain of f(x)= 3^x+2 is (-infinity, infinity)

OpenStudy (anonymous):

when you mean restrictions does that mean it equals zero?

jimthompson5910 (jim_thompson5910):

the same logic/reasoning applies to the other answer choices as well

jimthompson5910 (jim_thompson5910):

well by "restrictions" i mean values of x that are not allowed for example, the function \(\Large f(x) = \frac{1}{x-1}\) has the restriction that \(\Large x \neq 1\) because x = 1 makes the denominator 0 (and you can't divide by zero)

OpenStudy (anonymous):

ok and so when there are no x value restrictions the domain will always be (-infintity, infinity)?

jimthompson5910 (jim_thompson5910):

exactly

jimthompson5910 (jim_thompson5910):

so far, all choices are the answer but...we've yet to look at the range

OpenStudy (anonymous):

right and the range is all the output values

jimthompson5910 (jim_thompson5910):

all the possible output values, yes

jimthompson5910 (jim_thompson5910):

what is the range of the first answer choice?

OpenStudy (anonymous):

(2, infinity) since there is a positive 2 meaning as 2 approached infinity

jimthompson5910 (jim_thompson5910):

correct, the range of 3^x is (0,infinity) the +2 shifts up the range 2 units, so the 0 goes to 2

jimthompson5910 (jim_thompson5910):

how about choice B?

OpenStudy (anonymous):

(-2, infinity) since it shifts down 2 units

jimthompson5910 (jim_thompson5910):

very good

OpenStudy (anonymous):

C would be (2, infinity) and D. would be (-2,infinity). there fore the answers are A and C

jimthompson5910 (jim_thompson5910):

A and C are correct

jimthompson5910 (jim_thompson5910):

this is of course assuming the two functions for A and C are \[\Large f(x) = 3^x + 2\] and \[\Large f(x) = 0.25^x + 2\]

OpenStudy (anonymous):

now say you have an x value restriction would it just have an number instead of -infinity like say (1,infinity)? and yes that is what the two functions were

jimthompson5910 (jim_thompson5910):

what do you mean by your first question?

OpenStudy (anonymous):

like how you said 1/x-1 you cannot have x=1 b/c the denominator equals zero you cannot divide by zero so what would the domain be?

jimthompson5910 (jim_thompson5910):

oh if you have a restriction that x cannot equal 1, but it can equal any other number, then you basically start out with (-infintity, infinity) then you "poke a hole" at x = 1, aka you remove 1, to get this (-infintity, 1) U (1, infinity)

jimthompson5910 (jim_thompson5910):

Saying (-infintity, 1) U (1, infinity) is another way of saying "you can pick any number you want that's either less than 1 or greater than 1. You cannot choose 1 itself"

OpenStudy (anonymous):

ok thanks that makes a lot more since. i have a few quick questions to ask if you dont mind they are true and false i will tell you what i think it is i just want to know if i am right

OpenStudy (anonymous):

sense*

jimthompson5910 (jim_thompson5910):

sure, go ahead

OpenStudy (anonymous):

 i believe this is true